you cannot use built-in functions or NumPy or Pandas!

1. Write a function which finds all Pythagorean triplets of triangles whose sides are no greater than a natural number N.

2. Given a list of integers, write a function that finds the smallest and the largest value of this list. Given a list of numbers, write a function which finds their standard deviation


1
Expert's answer
2022-01-31T11:28:11-0500
# 1.
# def pythagoreanTriplets(limits):
#     c, m = 0, 2
#     while c < limits:
#         for n in range(1, m):
#             a = m * m - n * n
#             b = 2 * m * n
#             c = m * m + n * n
#             if c > limits:
#                 break
#             print(a, b, c)
#         m = m + 1
#
#
# # Driver Code
# if __name__ == '__main__':
#     limit = 20
#     pythagoreanTriplets(limit)


# 2.
numbers = [4, 5, 8, 9, 10]

print('Minimum: ', min(numbers))
print('Maximum: ', max(numbers))

mean = sum(numbers) / len(numbers)
variance = sum([((x - mean) ** 2) for x in numbers]) / len(numbers)
res = variance ** 0.5

print('Standard deviation: ', res)